### v1

@@ 5 Experiments
We conducted numerical experiments to evaluate the proposed filters and validate their theoretical properties. In Section˜5.1, we compare our methods against a wide range of time-varying and steady-state filters, assessing estimation accuracy, closed-loop performance in a trajectory-tracking task, and computational efficiency. We also examine how the ambiguity radii and the accuracy of the nominal noise distributions affect estimation performance. In Section˜5.2.2, we empirically verify the theoretical results of Section˜4.888All experiments were performed on a laptop equipped with an Intel Core Ultra 7 155H @ 3.80 GHz and 32 GB of RAM. Source code is available at https://github.com/jangminhyuk/DRKF2025.
@@ 5.1 Comparison of Robust State Estimators
We evaluate both the time-varying and steady-state DRKFs against the following benchmark state estimators:
@@
    DRKF with Wasserstein ambiguity on the joint law of $(x_{t},y_{t})$ [16]
@@
    DRKF with bicausal Wasserstein ambiguity [19]
@@
    Time-varying DRKF with Wasserstein ambiguity on the prior state and measurement noise [21, Theorem 1]
@@
    Time-invariant version of (G) with stationary nominals [21, Section 4]
@@
We use the same symbol $\theta$ for the robustness parameter in all filters, although its role and scale differ across methods. For (C) and (D), $\theta$ is the risk-sensitivity coefficient, with (D) using $\theta$ as the risk-sensitivity parameter. For (C), we select $\theta$ up to the largest value for which the risk-sensitive Riccati mapping remains well defined [35, 36], i.e., $((\Sigma_{x,t}^{-})^{-1}-\theta D^{\top}D)$ remains positive definite with $D=I_{n_{x}}$. For (E) and (F), $\theta$ specifies the Wasserstein radius, and for (F), we follow [19] and cap the number of internal iterations at 20. For (G) and (H), we set $\theta_{x}=\theta_{v}=\theta$. In our proposed formulations, we set $\theta_{x_{0}}=\theta_{w}=\theta_{v}=\theta$ for (I), and $\theta_{w}=\theta_{v}=\theta$ for (J). To ensure the fairness, we sweep each method’s admissible range of $\theta$ and report performance at its best tuning.
@@ 5.1.1 Estimation Accuracy
We begin with a linear time-invariant system (no control input) subject to inaccuracies in both the process and measurement noise distributions:
[…]
We consider two noise models: $(i)$ Gaussian, with $x_{0}\sim\mathcal{N}(0,0.01I_{4})$, $w_{t}\sim\mathcal{N}(0,0.01I_{4})$, $v_{t}\sim\mathcal{N}(0,0.01I_{2})$, and $(ii)$ U-Quadratic, with $x_{0}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $w_{t}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $v_{t}\sim\mathcal{UQ}([-0.1,0.1]^{2})$. The nominal covariances $\hat{\Sigma}_{w}$ and $\hat{\Sigma}_{v}$ are learned from only 10 input-output samples using the expectation-maximization (EM) method in [19]. This induces a significant distributional mismatch, highlighting the challenge of learning accurate nominal models from limited data. The estimation horizon is set to 50 time steps.
Fig.˜1 illustrates the effect of $\theta\in[10^{-2},10^{1}]$ on the average regret MSE for each filter under Gaussian and U-Quadratic noise, averaged over 20 runs. The regret MSE is defined as the difference between the filter’s MSE and that of the KF using the true noise distributions, which serves as the MMSE-optimal benchmark. The results show that both our time-varying and steady-state DRKFs consistently achieve the lowest tuned regret and remain competitive across a broad range of $\theta$. In particular, (I) achieves performance closest to the optimal estimator under both poorly estimated Gaussian and non-Gaussian noise.
In comparison, (F) exhibits numerical instability and performance discontinuities across $\theta$, even when the internal iteration limit is increased to 50. The risk-sensitive filters (C) and (D) become overly conservative for large $\theta$, leading to degraded performance. These results highlight the limitations of existing methods, whereas our DRKFs maintain stable performance and a favorable robustness-accuracy tradeoff.
Fig.˜2 shows the average MSE as a function of the number of samples used to construct the nominal noise distributions. Each filter is optimally tuned over $\theta\in[10^{-2},10^{1}]$, and we vary the number of input-output samples used in the EM procedure so that small sample sizes correspond to less accurate nominal distributions. As the number of samples increases, the average MSE decreases for all filters. Across all sample sizes, our DRKF consistently achieves the lowest error, with a particularly clear advantage when only limited data are available.
@@
Fig.˜3 shows the average MSE regret of the steady-state DRKF (J) as a function of the ambiguity radii $\theta_{w}$ and $\theta_{v}$ over $10^{-2},10^{1}]$, averaged over 100 runs. Both parameters significantly influence performance: the regret is large when both radii are small, reflecting insufficient robustness against distributional mismatch. Increasing either $\theta_{w}$ or $\theta_{v}$ improves performance up to an optimal range, beyond which excessively large radii make the estimator overly conservative. This surface highlights the need to balance $\theta_{w}$ and $\theta_{v}$ in practice.
@@ 5.1.2 Trajectory Tracking Task
To assess the practical performance of our filters, we applied them to a 2D trajectory-tracking task, following the setup in [21, Section 5.2]:
where $\Delta t=0.2\,\text{s}$. The state $x_{t}=[p_{t}^{x},p_{t}^{y},v_{t}^{x},v_{t}^{y}]^{\top}$ collects the planar position and velocity, and the control input $u_{t}=[a_{t}^{x},a_{t}^{y}]^{\top}$ represents acceleration commands.
We consider two noise settings: $(i)$ Gaussian, with $x_{0}\sim\mathcal{N}(0,0.02I_{4})$, $w_{t}\sim\mathcal{N}(0,0.02I_{4})$, $v_{t}\sim\mathcal{N}(0,0.02I_{2})$, and $(ii)$ nonzero-mean U-Quadratic, with $x_{0}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $w_{t}\sim\mathcal{UQ}([-0.1,0.2]^{4})$, $v_{t}\sim\mathcal{UQ}([-0.2,0.1]^{2})$, simulated over $10\,\text{s}$. For the nominal distributions, we fit an EM model using only $1\,\text{s}$ of input-output data, yielding noticeably inaccurate nominal statistics and reflecting the practical challenge of estimating noise models from limited observations.
Figure 5: 2D trajectories averaged over 200 simulations under Gaussian noise. Each filter uses its optimal parameter $\theta\in[10^{-2},10^{1}]$ minimizing the MPC cost. The black dashed line denotes the desired trajectory; colored curves show the mean, and shaded tubes indicate $\pm 1$ standard deviation.
@@
All methods use the same observer-based MPC controller (e.g., [40]), with the current state estimate $\bar{x}_{t}$ used for prediction. The controller solves a finite-horizon tracking problem with horizon $10$ and weights $Q=\mathrm{diag}([10,10,1,1])$ and $R=0.1I_{2}$, where $Q$ penalizes tracking error and $R$ penalizes control effort. The first control input $u_{t}$ is applied, and the optimization is repeated at the next step using $\bar{x}_{t+1}$. The MPC model, horizon, and weights are identical across all methods; the only component that changes is the state estimator, where the Kalman filter is replaced with each candidate estimator.
Fig.˜5 displays the mean and dispersion of the closed-loop trajectories obtained with the best-tuned radius $\theta\in[10^{-2},10^{1}]$ for each filter. The standard KFs (A) and (B) produce wider uncertainty tubes, indicating higher estimation variance. The BCOT-based DRKF (F) reduces dispersion but introduces greater bias, especially during sharp turns. The remaining DRKFs–(E), (G), and (H)–exhibit accurate tracking with modest dispersion. Our time-varying and steady-state DRKFs deliver the most reliable performance, achieving the lowest dispersion and bias across runs.
Although mean trajectories appear similar across DRKFs, the required control effort differs substantially. Table˜1 reports the time- and run-averaged acceleration magnitude $\|u_{t}\|_{2}$, which shows that our DRKFs achieve lower control energy among all methods.
5.1.3 Computation Time
The online computational cost of all filters is similar to that of the classical KF, since each estimator computes its gains offline. A key advantage of the steady-state DRKF is that it requires solving only a single SDP offline, whereas time-varying DRKFs must solve $T$ sequential SDPs.
Figure 6: Average offline computation time as a function of the horizon length $T$ for a system with $n_{x}=4$, $n_{y}=2$, averaged over 10 runs.
@@ 5.1.3 Computation Time
Fig.˜6 compares the offline computation time of (E)(J) as a function of the horizon length $T$. We use the same setup as in Section˜5.1.1 with $\theta=1.0$ and Gaussian noise. Time-varying DRKFs exhibit linear growth in computation time with respect to $T$, with (F) being particularly expensive. In contrast, the steady-state DRKFs (G) and (J) maintain constant runtime regardless of $T$, demonstrating excellent scalability for long-horizon tasks.
@@ 5.2 Validation of Theoretical Properties
5.2.1 DRKF Sandwich Properties
Theorem˜2 states that, at each time step, the DRKF prior and posterior covariances lie between those of two KFs run with extremal noise levels from (17). We illustrate this sandwich property on a concrete 2D example. Consider the system
$A=0.95\begin{bmatrix}\cos(\pi/8)&-\sin(\pi/8)\\ \sin(\pi/8)&\cos(\pi/8)\end{bmatrix},\quad C=\begin{bmatrix}1.0&0.1\\ 0.1&1.0\end{bmatrix}$,
with horizon $T=20$ and ambiguity radii $\theta_{w}=\theta_{v}=\theta_{x_{0}}=0.1$. The initial prior covariance is
$\Sigma_{x,0}^{-}=\begin{bmatrix}0.2&0.05\\ 0.05&0.1\end{bmatrix}$.
We use time-varying nominal covariances of the form $\hat{\Sigma}_{w,t}=a_{t}\hat{\Sigma}_{w}$, $\hat{\Sigma}_{v,t}=a_{t}\hat{\Sigma}_{v}$, constructed by scaling the base matrices
@@ 5.2.1 DRKF Sandwich Properties
These covariances are used at each step $t$ to generate the lower-bound KF (LOW-KF), upper-bound KF (HIGH-KF), and DRKF posterior covariance trajectories. To visualize the uncertainty, we use the $95\%$ confidence ellipse $\mathcal{E}(\Sigma)\coloneqq\{x\in\mathbb{R}^{2}:x^{\top}\Sigma^{-1}x\leq\chi^{2}_{2,0.95}\}$ with $\chi^{2}_{2,0.95}=5.991$. This region contains $95\%$ of a zero-mean Gaussian with covariance $\Sigma$. By Theorem˜2, $\mathcal{E}(\underline{\Sigma}_{x,t})\subseteq\mathcal{E}(\Sigma_{x,t})\subseteq\mathcal{E}(\overline{\Sigma}_{x,t}),$ so the DRKF posterior ellipse must lie between the LOW-KF and HIGH-KF ellipses at every time step.
As shown in Figs.˜7 and 8, the ellipses at representative time steps confirm the predicted nesting. The DRKF posterior covariance is consistently sandwiched between the LOW-KF and HIGH-KF bounds, providing a clear and interpretable robustness envelope throughout the horizon.
5.2.2 Convergence Properties
To evaluate the convergence behavior of the proposed filter, we consider the linear system from [37]:
$A=\begin{bmatrix}0.1&1\\ 1&-1\end{bmatrix},\;C=\begin{bmatrix}1&-1\end{bmatrix}$.
Figure 9: Relative difference between $\operatorname{Tr}[\Sigma_{x,t}]$ and $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ over time.
We set the nominal covariances to $\hat{\Sigma}_{w}=I_{2}$ and $\hat{\Sigma}_{v}=1$. Following the procedure in Section˜4.2, we first compute $\phi_{N}$ with $N=8$ and $q=20$, and set $\phi_{v}=\phi_{x}=\phi_{N}/2$. We then choose the largest pair $(\theta_{v},\theta_{w})$ permitted by Theorem˜3.
To quantify convergence of the time-varying DRKF toward the steady-state solution, we track the relative error $|(\operatorname{Tr}[\Sigma_{x,\infty}^{*}]-\operatorname{Tr}[\Sigma_{x,t}])/\operatorname{Tr}[\Sigma_{x,\infty}^{*}]|$. As shown in Fig.˜9, this quantity decays approximately linearly on a logarithmic scale, demonstrating exponential convergence of the time-varying DRKF to the steady-state DRKF.
6 Conclusions and Future Work
This work presented a noise-centric formulation of DRKF in which uncertainty in the process and measurement noises is modeled directly via Wasserstein ambiguity sets. In contrast to prior DRKF formulations, the proposed approach preserves dynamic consistency and yields tractable spectral bounds on the least-favorable covariances. The resulting recursion retains the analytical structure of the classical KF while providing robustness to distributional errors. We established the existence, uniqueness, and convergence of the steady-state DRKF and proved that it is asymptotically minimax optimal with respect to the worst-case mean-square error. Notably, the steady-state solution is obtained from a single stationary SDP, yielding a constant-gain filter with Kalman-level online complexity and strong robustness guarantees. Empirical studies corroborate the theory, demonstrating improved estimation accuracy, tighter and more reliable uncertainty quantification, and superior closed-loop performance in control tasks.
Future directions include extending the framework to nonlinear systems and developing adaptive mechanisms for online calibration of the ambiguity set radius. It is also of interest to relax some of the structural assumptions underlying our convergence analysis (e.g., independence and range conditions on the process noise), to incorporate simultaneous model uncertainty in the system matrices, and to explore temporally coupled ambiguity sets that better capture long-range noise correlations while preserving computational tractability.
@@ Appendix A Eigenvalue Lower Bounds for DRSE Maximizers
For $t=0$, the DRSE formulation (4) admits a maximizer pair $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ such that
$\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}.$ (33)
For $t>0$, the DRSE formulation (5) admits a maximizer pair $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ such that
$\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}.$ (34)
@@
For $t=0$, (33) is shown in [17, Theorem 3.1]. We now prove (34) by reformulating (5) in a form to which [17, Lemma A.3] applies. The objective of (5) is the trace of the posterior covariance matrix:
$\begin{split}\operatorname{Tr}[\Sigma_{x,t}]&=\operatorname{Tr}[\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C^{\top}(C\Sigma_{x,t}^{-}C^{\top}+\Sigma_{v,t})^{-1}C\Sigma_{x,t}^{-}]\\ &=\inf_{K_{t}}\operatorname{Tr}[(I_{n_{x}}-K_{t}C_{t})\Sigma_{x,t}^{-}(I_{n_{x}}-K_{t}C_{t})^{\top}+K_{t}\Sigma_{v,t}K_{t}^{\top}]\\ &=\inf_{K_{t}}(\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\Sigma_{x,t}^{-}\rangle+\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle)\end{split}$
where the last equality uses the inner product $\langle A,B\rangle:=\operatorname{Tr}[A^{\top}B]$. Given $\Sigma_{x,t-1}$, the deterministic part $A_{t}\Sigma_{x,t-1}A_{t}^{\top}$ is fixed, so $\Sigma_{x,t}^{-}=\bar{\Sigma}_{t}+\Sigma_{w,t-1}$, $\bar{\Sigma}_{t}\coloneqq A_{t}\Sigma_{x,t-1}A_{t}^{\top}\succeq 0$. Substituting this into the objective yields
$\begin{split}\operatorname{Tr}[\Sigma_{x,t}]&=\inf_{K_{t}}\bigg(\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\Sigma_{w,t-1}\rangle\\ &\qquad+\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\bar{\Sigma}_{t}\rangle+\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle\bigg)\end{split}$ (35)
First, we fix $\Sigma_{v,t}$ and view the objective as a function of $\Sigma_{w,t-1}$. From (35) this dependence can be written as
$\inf_{K_{t}}\Big(\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\Sigma_{w,t-1}\rangle+f_{1,t}(K_{t})\Big),$ (36)
where $f_{1,t}(K_{t})$ collects the remaining terms, and is convex and continuous in $K_{t}$. Define the convex set $\mathcal{C}\coloneqq\{I_{n_{x}}-K_{t}C_{t}:K_{t}\in\mathbb{R}^{n_{x}\times n_{y}}\}$. For each $L\in\mathcal{C}$, define
$f(L)\coloneqq\inf_{K_{t}:I_{n_{x}}-K_{t}C_{t}=L}f_{1,t}(K_{t}).$
The function $f$ is convex and continuous since $f_{1,t}$ is quadratic in $K_{t}$ and we take the infimum over an affine subspace. Then (36) can be rewritten as
$\inf_{L\in\mathcal{C}}\big(\langle L^{\top}L,\Sigma_{w,t-1}\rangle+f(L)\big).$
Together with the Bures-Wasserstein constraint on $\Sigma_{w,t-1}$ in (5), this matches the setting of [17, Lemma A.3] with variable $\Sigma=\Sigma_{w,t-1}$. Hence there exists a maximizer $\Sigma_{w,t-1}^{*}$ such that $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}}$.
Next, we fix $\Sigma_{w,t-1}$ (and thus $\Sigma_{x,t}^{-}$) and view the objective as a function of $\Sigma_{v,t}$. From (35) the dependence on $\Sigma_{v,t}$ is
$\inf_{K_{t}}\Big(\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle+f_{2,t}(K_{t})\Big),$
where $f_{2,t}(K_{t})$ collects the remaining terms and is convex and continuous in $K_{t}$. This is exactly of the form in [17, Lemma A.3] with variable $\Sigma=\Sigma_{v,t}$. Applying [17, Lemma A.3] again yields a maximizer $\Sigma_{v,t}^{*}$ such that $\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$.
By choosing such partial maximizers when forming a joint maximizer of (5), we obtain (34), completing the proof.

### v3

@@ 5 Experiments
We evaluate the proposed DRKFs through numerical experiments that assess estimation accuracy, closed-loop performance, and computational efficiency, and that empirically validate the theoretical results in Section˜4.101010All experiments were performed on a laptop equipped with an Intel Core Ultra 7 155H @ 3.80 GHz and 32 GB of RAM. Source code is available at https://github.com/jangminhyuk/DRKF2025.
@@ 5.1 Comparison of Robust State Estimators
We begin by comparing the proposed time-varying and steady-state DRKFs with the following baseline estimators:
@@
    DRKF with joint Wasserstein ambiguity on $(x_{t},y_{t})$ [26]
@@
    DRKF with bicausal Wasserstein ambiguity [8] [^11] Following [8], we cap the number of internal iterations at 20.
@@
    Time-varying DRKF with Wasserstein ambiguity on the prior state and measurement noise [11, Thm. 1]
@@
    Time-invariant version of (G) with stationary nominals [11, Sect. 4]
@@
We use the same symbol $\theta$ to denote the robustness parameter across all methods, although its interpretation differs. For risk-sensitive filters (C) and (D), $\theta$ is the risk parameter (with opposite signs for risk-averse/seeking variants), chosen within the stability range [16]. For Wasserstein-based methods, $\theta$ denotes the ambiguity radius. For (G) and (H), we set $\theta_{x_{0}}=\theta_{v}=\theta$. In our proposed formulations, we set $\theta_{x_{0}}=\theta_{w}=\theta_{v}=\theta$ for (I), and $\theta_{w}=\theta_{v}=\theta$ for (J). To ensure fairness, we sweep each method over its admissible range of $\theta$ and report performance at the best tuned value. [^12] In practice, DRKF radii $\theta$ can be selected from data, following standard approaches in the DRO literature, such as cross-validation or bootstrapping (e.g., [21, 4]).
@@ 5.1.1 Estimation Accuracy
We first consider a linear time-invariant system, subject to inaccuracies in both the process and measurement noise distributions:
[…]
We test two noise models: $(i)$ Gaussian, with $x_{0}\sim\mathcal{N}(0,0.01I_{4})$, $w_{t}\sim\mathcal{N}(0,0.01I_{4})$, $v_{t}\sim\mathcal{N}(0,0.01I_{2})$, and $(ii)$ U-Quadratic, with $x_{0}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $w_{t}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $v_{t}\sim\mathcal{UQ}([-0.1,0.1]^{2})$. Nominal covariances are learned from only 10 measurement samples using the expectation-maximization (EM) method in [8], creating a significant mismatch between nominal and true noise statistics. The estimation horizon is $T=50$.
Fig.˜1 shows the effect of $\theta\in[10^{-2},10^{1}]$ on the average regret MSE under both noise models, averaged over 2020 runs. The regret MSE is defined as the difference between the filter’s MSE and that of the KF using the true noise distributions. The results show that both proposed DRKFs consistently achieve the lowest tuned regret and remain competitive across a wide range of $\theta$. In particular, (I) achieves performance closest to the optimal estimator under both poorly estimated Gaussian and non-Gaussian noise.
In contrast, the BCOT-based DRKF (F) exhibits numerical instability and discontinuous performance across $\theta$, even when increasing the internal iteration limit to 50. The risk-sensitive filters (C) and (D) become overly conservative for large $\theta$, leading to degraded accuracy.
@@
Figure 5: 2D trajectories averaged over 200 simulations. Each filter uses its optimal $\theta\in[10^{-2},10^{1}]$ minimizing the MPC cost. The dashed black line is the desired trajectory; colored curves show the mean, and shaded tubes indicate $\pm 1$ standard deviation.
Fig.˜2 further examines performance as a function of the number of samples used to estimate the nominal noise statistics. [^13] Each filter is optimally tuned over $\theta\in[10^{-2},10^{1}]$, and we vary the number of input-output samples used in the EM procedure so that small sample sizes correspond to less accurate nominal distributions. Across all sample sizes, the proposed DRKFs achieve the lowest MSE, with particularly clear advantages in the low-data regime.
Fig.˜3 illustrates the sensitivity of the steady-state DRKF to the ambiguity radii $(\theta_{w},\theta_{v})$. Insufficient robustness (small radii) leads to large regret, while overly large radii induce excessive conservatism, highlighting the need to balance $\theta_{w}$ and $\theta_{v}$.
@@ 5.1.2 Trajectory Tracking Task
We next evaluate closed-loop performance on a 2D trajectory-tracking task, following the setup in [11, Sect. 5.2]:
where $\Delta t=0.2\,\text{s}$. The state $x_{t}=[p_{t}^{x},p_{t}^{y},v_{t}^{x},v_{t}^{y}]^{\top}$ collects the planar position and velocity, and the control input $u_{t}=[a_{t}^{x},a_{t}^{y}]^{\top}$ represents acceleration commands. All methods use the same observer-based MPC controller, with the current state estimate $\bar{x}_{t}$ provided by each filter. The controller solves a finite-horizon tracking problem with horizon $10$ and weights $Q=\mathrm{diag}([10,10,1,1])$ and $R=0.1I_{2}$, where $Q$ penalizes tracking error and $R$ penalizes control effort. The first control input $u_{t}$ is applied, and the optimization is repeated at the next step using $\bar{x}_{t+1}$.
The true noise distributions are $x_{0}\sim\mathcal{N}(0,0.02I_{4})$, $w_{t}\sim\mathcal{N}(0,0.02I_{4})$, and $v_{t}\sim\mathcal{N}(0,0.02I_{2})$, simulated over $10\,\text{s}$. For the nominal distributions, we fit an EM model using only $1\,\text{s}$ of input-output data, yielding noticeably inaccurate nominal statistics and reflecting the practical challenge of estimating noise models from limited observations.
Figure 6: Average offline computation time as a function of the horizon length $T$ for $n_{x}=4$, $n_{y}=2$, averaged over 10 runs.
@@
Fig.˜5 displays the mean trajectories and uncertainty envelopes obtained with the best-tuned radius $\theta\in[10^{-2},10^{1}]$ for each filter. The standard KFs (A) and (B) produce wider uncertainty tubes, indicating higher estimation variance. The BCOT-based DRKF (F) reduces dispersion but introduces greater bias, especially during sharp turns. Other DRKFs–(E), (G), and (H)–achieve improved tracking, while the proposed time-varying and steady-state DRKFs exhibit the lowest dispersion and bias.
Although mean trajectories are similar across DRKFs, control effort differs substantially. As shown in Table˜1, the proposed DRKFs require the lowest average control energy among all methods.
5.1.3 Computation Time
Fig.˜6 compares the offline computation time of (E)(J) as a function of the horizon length TT. [^14] We use the same setup as in Section 5.1.1 with $\theta=1.0$ and Gaussian noise. Time-varying DRKFs exhibit linear growth in offline computation time with respect to $T$, with (F) being particularly expensive. In contrast, the steady-state DRKFs (H) and (J) maintain constant runtime, as they require solving only a single SDP offline. This demonstrates the scalability of the proposed steady-state DRKF for long-horizon applications.
@@ 5.2 Validation of Theoretical Properties
5.2.1 Covariance Sandwich Property
We illustrate the covariance sandwich property of Theorem˜2 using a 2D example with
$A=0.95\begin{bmatrix}\cos(\pi/8)&-\sin(\pi/8)\\ \sin(\pi/8)&\cos(\pi/8)\end{bmatrix},\qquad C=\begin{bmatrix}1.0&0.1\\ 0.1&1.0\end{bmatrix}$,
horizon $T=20$, and ambiguity radii $\theta_{w}=\theta_{v}=\theta_{x_{0}}=0.1$. The initial prior covariance is $\Sigma_{x,0}^{-}=\begin{bmatrix}0.2&0.05\\ 0.05&0.1\end{bmatrix}$. We use time-varying nominal covariances of the form $\hat{\Sigma}_{w,t}=a_{t}\hat{\Sigma}_{w}$, $\hat{\Sigma}_{v,t}=a_{t}\hat{\Sigma}_{v}$, constructed by scaling the base matrices
@@ 5.2.1 Covariance Sandwich Property
These covariances are used at each step $t$ to generate the lower-bound KF (LOW-KF), upper-bound KF (HIGH-KF), and DRKF posterior covariance trajectories. To visualize the uncertainty, we use the $95\%$ confidence ellipse $\mathcal{E}(\Sigma)\coloneqq\{x\in\real{2}:x^{\top}\Sigma^{-1}x\leq\chi^{2}_{2,0.95}\}$ with $\chi^{2}_{2,0.95}=5.991$. This region contains $95\%$ of a zero-mean Gaussian with covariance $\Sigma$. By Theorem˜2, $\mathcal{E}(\underline{\Sigma}_{x,t})\subseteq\mathcal{E}(\Sigma_{x,t})\subseteq\mathcal{E}(\overline{\Sigma}_{x,t})$.
As shown in Fig.˜7, the ellipses at representative time steps confirm the predicted nesting, providing a clear and interpretable robustness envelope.
5.2.2 Convergence Property
Finally, we examine convergence of the time-varying DRKF to its steady-state counterpart using the example from [34]: $A=\begin{bmatrix}0.1&1\\ 1&-1\end{bmatrix},\quad C=\begin{bmatrix}1&-1\end{bmatrix}$. [^15] We set the nominal covariances to $\hat{\Sigma}_{w}=I_{2}$ and $\hat{\Sigma}_{v}={1}$. and the ambiguity set radii $\theta_{x_{0}}=\theta_{w}=\theta_{v}=0.1$. Fig.˜9 shows that the relative error between $\operatorname{Tr}[\Sigma_{x,t}]$ and $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ decays approximately linearly on a logarithmic scale, confirming exponential convergence, consistent with the theoretical analysis.
Figure 9: Relative difference between $\operatorname{Tr}[\Sigma_{x,t}]$ and $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$.
6 Conclusions
We developed a DRKF based on a noise-centric Wasserstein ambiguity model. The proposed formulation preserves the structure of the classical KF, admits explicit spectral bounds, and ensures dynamic consistency under distributional uncertainty. We established existence, uniqueness, and convergence of the steady-state solution, proved asymptotic minimax optimality, and showed that the steady-state filter is obtained from a single stationary SDP with Kalman-level online complexity. Future directions include extending the framework to nonlinear systems, developing adaptive mechanisms for online calibration of the ambiguity set radius, incorporating model uncertainty in the system matrices, and exploring temporally coupled ambiguity sets.
@@ Appendix A Eigenvalue Lower Bounds for DRSE Maximizers
Consider the problems (4) and (5) with the explicit eigenvalue lower-bound constraints removed. Then:
For $t=0$, (4) admits a maximizer $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ such that
$\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\quad\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}.$ (31)
For $t>0$, (5) admits a maximizer $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ such that
$\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\quad\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}.$ (32)
@@
For $t=0$, (31) is shown in [23, Thm. 3.1]. We prove (32) for $t>0$. Fix $\Sigma_{x,t-1}$, so that $\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}$. Using the standard identity $\operatorname{Tr}[\Sigma_{x,t}]=\inf_{K_{t}}\operatorname{Tr}\left[(I-K_{t}C_{t})\Sigma_{x,t}^{-}(I-K_{t}C_{t})^{\top}+K_{t}\Sigma_{v,t}K_{t}^{\top}\right]$, the objective can be written as
$\operatorname{Tr}[\Sigma_{x,t}]=\inf_{K_{t}}\Big(\langle(I-K_{t}C_{t})^{\top}(I-K_{t}C_{t}),\,\Sigma_{w,t-1}\rangle+f_{t}(K_{t},\Sigma_{v,t})\Big),$
where $\langle A,B\rangle\coloneqq\operatorname{Tr}[A^{\top}B]$ and $f_{t}$ collects the remaining terms (independent of $\Sigma_{w,t-1}$) and is convex and continuous in $K_{t}$.
Let $(\bar{\Sigma}_{w,t-1},\bar{\Sigma}_{v,t})$ be a maximizer of (5) with the explicit eigenvalue bounds removed, which exists since the feasible set (Bures–Wasserstein balls intersected with $\mathbb{S}^{\cdot}_{+}$) is compact and the objective is continuous. Fixing $\Sigma_{v,t}=\bar{\Sigma}_{v,t}$, the maximization over $\Sigma_{w,t-1}$ matches the setting of [23, Lemma A.3], hence it admits a maximizer $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}}$. Moreover, because $(\bar{\Sigma}_{w,t-1},\bar{\Sigma}_{v,t})$ is globally optimal, the fixed-$\Sigma_{v,t}$ subproblem achieves the same optimal value, so $(\Sigma_{w,t-1}^{*},\bar{\Sigma}_{v,t})$ is also a global maximizer.
Now fix $\Sigma_{w,t-1}=\Sigma_{w,t-1}^{*}$, which in turn fixes $\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}^{*}$. Then, the dependence on $\Sigma_{v,t}$ is of the form $\inf_{K_{t}}(\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle+\tilde{f}_{t}(K_{t}))$, so applying [23, Lemma A.3] again yields a maximizer $\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$, and $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ remains globally optimal. This proves (32).